16. Oktober 2003. U. Schoenwaelder; http://www.math.rwth-aachen.de/∼Ulrich.Schoenwaelder HB = Hochschulbibl. RWTH, HBZ = http://www.hbz-nrw.de/ (HBZ-CD-ROM Online), MB = Mathematikbibl., DB = Didaktikbibl. (Winter), EB = Erziehungswiss. Bibl., FH = Bibl. Fachhochschule Aachen, FL = Fernleihe, IB = Informatikbibliothek, IB Nr. = Institutsbibliothek Nr., LB = HB– Lehrbuchsammlung, LS = HB–Lesesaal

Literatur zu figurierten Zahlen

[1] J. H. Conway and R. K. Guy. The Book of Numbers. Springer–Verlag, 1996. MB: 17923. S. 17: How numbers are written; S. 30: Square numbers; S. 33: Triangular numbers; polygonal numbers; tetrahedral numbers; sums of cubes; S. 68: Pascal’s triangle; S. 91: Bell numbers, Stirlilng numbers, Catalan numbers, Bernoulli numbers, Fibonacci numbers (sumflower) [see also S. 202]; S. 127: Primes; S. 146: Sums of two squares; S. 152: Farey fractions and Ford circles; S. 157: Fractions cycle into decimals; S. 171: Pythagorean fractions; S. 176: Continued fractions; S. 181: Geometric problems and algebraic numbers; S. 211: Complex numbers; Gaussian primes; Eisenstein primes; S. 230: Hamilton’s quaternions; S. 237: Some transcendental numbers: π, Liouville’s number, e, ...; S. 265: Infinite and infinitesimal numbers; S. 283: Surreal numbers; games.

[2] Stefanie Krivsky et al. MathePrisma. http://www.MathePrisma.uni-wuppertal.de, Gesehen Oktober 2001. Interview in DIE ZEIT Nr. 41, 4. Oktober 2001, Seite 80 Chancen. Enthält den Modul ”Quadratzahlen“ mit Dreieckszahlen, Pythagoras.

[3] J. A. Ewell. On sums of triangular numbers and sums of squares. Amer. Math. Monthly, 99(8):752–757, 1992. MB: Z 42. Triangular number = n(n + 1)/2.

[4] T. L. Heath. Diophantus of Alexandria. A Study in the History of Greek Algebra. Dover, 1964. Originally 1910. P. 127: Polygonal Numbers.

[5] Detlef Laugwitz. On Polygonal Numbers. TU Darmstadt/FB Mathematik Preprint 1807 (5 S.), 1996. HBZ: 466 PB.

[6] Mason, Graham, Pimm, and Gowar. Routes to / Rootes of Algebra. Open University, 1985. Problems on figural numbers require students to see, say, record and test a pattern.

[7] Mathe-Welt. Mathe-Welt: Quadratzahlen, Dreieckszahlen, Fibonacci-Zahlen. mathematik lehren 87, 31–33, April 1998.

[8] Michael Neubrand and Manfred Möller. Einführung in die Arithmetik: ein Arbeitsbuch für Studierende des Lehramts der Primarstufe. Franzbecker, 1990, ^{2}1992. FL: UB Essen TOI 1158+1. ISBN 3-88120-193-9. Kap. 4: Zahlen und Muster (Dreieckszahlen, Quadratzahlen, Sechseckzahlen, Kubikzahlen).

[9] Luis Radford. The roles of geometry and arithmetic in the development of algebra: historical remarks from a didactic perspective. In Nadine Bednarz, Carolyn Kieran, and Lesley Lee, editors, Approaches to Algebra: Perspectives for Research and Teaching, Mathematics Education Library 18, pages 39–53. Kluwer, 1996. FL: UB Duisburg 01 TOP 1111. ISBN 0-7923-4145-7. Polygonal numbers emerge in a philosophical context of classification of numbers (square, cube, ..) which dates back to the era of the first Pythagoreans.

[10] D. D. Spalt. Der induktive Diskurs - Lakatos zum Induktionsproblem in der Mathematik. In M. Glatfeld, editor, Überlegungen zum Induktionsbegriff – unter fachdidaktischer Hinsicht, pages 101–159? Peter Lang, 1987. Ka 5553-287+1. S. 153: Figurierte Zahlen und Induktion.

[11] M. L. D’Ooge (Transl.), F. E. Robbins, and L. C. Karpinski (Eds.). Nicomachus of Gerasa. Introduction to Arithmetic. With Studies in Greek Arithmetic. Macmillan, 1926. Square numbers, sum of consecutive triangle numbers.

[12] Heinrich Winter. Quadrat und Zahl; ästhetische Erfahrungen im Mathematikunterricht. In Thomas Weth, editor, Kreativität, mathematik lehren 106, pages 19–22, 39–41. Friedrich, Juni 2001. Das Aufdecken von Wechselbeziehungen zwischen Gestalt und Zahl ist eine der kreativsten Formen des Mathematiklernens, weil dabei ästhetische Momente wirksam sind. Quadratzahlen, der euklische Algorithmus, Fibonacci-Zahlen, Bruchrechnung mit Quadraten.